Which of these angles is not coterminal with the others?
11pie/6
17pie/6
-7pie/6
-19pie/6

Answer by stanbon(75887)

Which of these angles is not coterminal with the others?
11pie/6 = 1 5/6 pi
17pie/6 = 2 5/6 pi = 5/6 pi
-7pie/6 = -1 1/6 pi
-19pie/6 = -3 1/6 pi = -1 1/6 pi
---------
Ans: 3rd and 4th are coterminal
=======================
Cheers,
Stan H.